OpenStudy (anonymous):

Does anyone know what 1/3+1/6 equals ?

OpenStudy (anonymous):

2/6 + 1/6

OpenStudy (anonymous):

aka, 3/6 or 1/2

OpenStudy (anonymous):

? Can you explain it ?

OpenStudy (anonymous):

Dont you have to find the GCF ?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

1/3+1/6 we need to make the denominators equal the easiest way to do that is by multiplying the fraction 1/3 by 2 so now we have 2/6+1/6 then we just add normally

OpenStudy (anonymous):

First, find the common denominator (the number on the bottom of the fraction). In this case, the common denominator of 3 and 6 is 6, because 3x2 equals 6, and 6x1 can also equal 6. 6 is your common denominator, or GCF. 1/3 = something/6 How do you find the top number? Well, to get 6, you had to multiply 3 by 2. To get your top number, multiplt 1 by 2 as well. So you have: 1/3= 2/6. Since we only had to multiply 1/6 by 1 to get the common denomintor, it virtually stays the same. Now we add 1/6 and 2/6 to get 3/6, which can be simplified to 1/2.

OpenStudy (anonymous):

you can have this question jesssielynnn33

OpenStudy (anonymous):

I'm sorry Micheal.3545 I didn't mean to be rude! Just trying to help

OpenStudy (anonymous):

Thanks SO much for explaining it to me.

OpenStudy (anonymous):

Bubblegum, if you are still confused, this video may help. http://www.khanacademy.org/video/adding-and-subtracting-fractions?playlist=Pre-algebra

OpenStudy (anonymous):

Thank You